A rectangular parking lot is 100 ft longer than it is wide. Determine the dimensions of the parking lot if it measures 500 ft diagonally.
Found 2 solutions by midwood_trail, csteadman:
Answer by midwood_trail(310)

A rectangular parking lot is 100 ft longer than it is wide. Determine the dimensions of the parking lot if it measures 500 ft diagonally.
length = x + 100 feet
width = x feet
I will set it up and you can finish.
The diagonal length = 500 feet
Use the Pythegorean Theorem.
(x)^2 + (x + 100)^2 = (500)^2
When you find x, plug your answer for x to find the length and width.
Can you take it from here?
If not, write back.

Answer by csteadman(2)